# Parametric Channel Estimation and Design for Active-RIS-Assisted Communications

### v1

@@ IV Adaptive Beam and Amplification Factor
4:for $\ell=2$ to $L$ do
5:  $\mathbf{y}_{\ell}=\sqrt{P_{p}}\mathbf{B}_{\ell}\mathbf{D}_{h}\mathbf{g}+\mathbf{v}_{\ell}+\mathbf{w}_{\ell}$
6: where $\mathbf{v}_{\ell}]_{i}=\mathbf{B}_{\ell}(i,:)\mathbf{D}_{h}\mathbf{v}_{i}$
7:  Compute closed-form MLE:
$\hat{\bm{\psi}}_{\ell}=\arg\max_{\bm{\psi}\in\Psi}\frac{\left|\mathbf{y}^{H}\mathbf{F}^{-1}\mathbf{B}\mathbf{D}_{h}\bm{a(\psi)}\right|^{2}}{\bm{{a}(\psi)}^{H}\mathbf{D}_{h}^{H}\mathbf{B}^{H}\mathbf{F}^{-1}\mathbf{B}\mathbf{D}_{h}\bm{{a}(\psi)}}$
8:  Obtain $\hat{\beta}_{\ell}$, $\hat{\omega}_{\ell}$ from (14) and (12) and set $\hat{\mathbf{g}}_{\ell}$ according to (3)
9:  Active RIS amplitude shaping: for $k=1,\ldots,N$
10: Compute $p_{k}^{(\mathrm{est})}=C\dfrac{\alpha_{k}}{\beta_{k}+\gamma_{k}}$
11:  if $\ell<L$ then
12:   Compute:
$\bar{\bm{\theta}}_{\ell}^{\mathrm{opt}}=\exp\left(-j\,\angle(\mathbf{D}_{h}\hat{\mathbf{g}}_{\ell})\right),\quad\bm{\phi}_{\ell}^{\star}=\mathbf{p}^{(\mathrm{est})}\odot\bar{\bm{\theta}}_{\ell}^{\mathrm{opt}}$
13:   Select next pilot via:
$\bm{\theta}_{\ell+1}=\arg\max_{\bm{\theta}\in\bm{\Theta}}\left|(\bm{\phi}_{\ell}^{\star})^{\mathrm{H}}\bm{\theta}\right|$
14:   Update:
$\bm{\phi}_{\mathrm{new}}=\mathbf{p}^{(\mathrm{est})}\odot\bm{\theta}_{\ell+1},\quad\mathbf{B}_{\ell+1}=\begin{bmatrix}\mathbf{B}_{\ell}\\ \bm{\phi}_{\mathrm{new}}^{\top}\end{bmatrix}$,
$\quad\bm{\Theta}\leftarrow\bm{\Theta}\setminus\{\bm{\theta}_{\ell+1}\}.$
15:   Transmit pilot $\ell+1$ and collect $\mathbf{y}_{\ell+1}$
16:  end if
17:end for
18:Output: $\hat{\mathbf{g}}=\hat{\mathbf{g}}_{L}$
Adaptive pilot selection via the closest-beam rule:
To refine the estimate with additional pilots, we choose the next phase-only codeword by maximizing its alignment with the current active vector:
Here $\bm{\theta}^{\text{next}}$ is the phase-only active RIS setting for the next pilot. The corresponding active RIS vector is
$\bm{\phi}_{\text{new}}\,=\,\bm{p}^{(\mathrm{est})}\,\odot\,\bm{\theta}^{\text{next}}.$ (22)
We append $\bm{\phi}_{\text{new}}$ as a new row to the configuration matrix $\mathbf{B}$, collect the corresponding measurement, and update $\widehat{\bm{g}}$ via the MLE step. This greedy closest-beam procedure repeats until the pilot budget is exhausted. After selecting a configuration from the codebook $\Theta$, it is removed from further consideration.
Figure 1: NMSE (Near-field)
Figure 2: Rate (Near-field)
Figure 3: NMSE (Far-field)
Figure 4: Rate (Far-field)
Remarks on complexity and applicability:
At each refinement, we evaluate the remaining phase-only codewords in the codebook. Let ${\Theta}_{\text{rem}}$ denote the unused set of RIS configuration after the two initial pilots. Each iteration computes $|{\Theta}_{\text{rem}}|$ inner products of length $N$, giving a per-iteration complexity of $\mathcal{O}\!\left(N\,|{\Theta}_{\text{rem}}|\right)$. With $L$ instances of the pilot and total codebook size $R$ (with $R\ll N^{2}$), the overall training cost scales as $\mathcal{O}(LNR)$, which is substantially lower than DFT-based exhaustive search. The procedure is agnostic to the propagation regime: only the steering parameterization $\bm{\psi}$ changes—angles in the far field, and angles plus distance in the near field—while the MLE and adaptive-pilot logic remain the same.
@@ V Simulation Results Analysis
We evaluate the proposed estimator’s performance through Monte Carlo simulations. The system operates at a carrier frequency of $28$ GHz with a $1$ MHz bandwidth and consists of a single-antenna base station (BS) and user equipment (UE) assisted by an active reconfigurable intelligent surface (ARIS). The active RIS is a uniform planar array of $32\times 32$ elements with half-wavelength spacing ($\Delta_{H}=\Delta_{V}=\lambda/2$), positioned $15$ m from the BS.
The users are randomly distributed with azimuth and elevation angles drawn from $\phi,\varphi\sim\mathcal{U}[-\pi/3,\pi/3]$. We distinguish between near- and far-field scenarios based on the Fraunhofer distance, $d_{\text{f}}=2D^{2}/\lambda$, where $D$ is the largest physical aperture dimension. For the near-field, the user distance $r$ is sampled from $\mathcal{U}[d_{\text{B}},d_{\text{f}}/10]$, where $d_{\text{B}}=2D$ denotes the Björnson distance [15]. For the far-field, the distance is drawn from $\mathcal{U}[d_{\text{f}},5d_{\text{f}}]$. The pilot transmission power is set $10$ dB above the data power.
To validate the effectiveness of the active architecture, the proposed scheme is benchmarked against a conventional passive RIS. The most crucial aspect of this comparison is maintaining a fair total power budget. For the passive RIS scenario, the entire power budget is dedicated to the BS transmission. On the other hand, for the active RIS scenario, $25\%$ of the total power is allocated to the RIS’s internal amplifiers. Consequently, the BS transmits power in the active RIS-assisted system is limited to the remaining $75\%$ of the total budget, ensuring an equitable comparison between the two architectures. $200$ mW is set for the total power. The presented results are averaged over $1000$ independent channels and noise realizations.
@@
For the case of the user being in the near field, the active RIS achieves a final NMSE that is considerably lower by several dB than the minimum NMSE achieved by the passive RIS structure. This behavior is being observed under both the near and far field approximations. In contrast, for the case of the user being in the far field, both these architectures portray almost equivalent behavior under both approximation models.
The rate curves in Figs. 4 and 4 represent the theoretical capacity limits as well as the pilots achievable rate. It consistently shows higher values for the active RIS compared to the passive RIS for all the pilots. Crucially, the operational rate of the active RIS closely approaches its capacity bound, whereas the passive RIS operates well below that of the active RIS, reflecting a much lower capacity ceiling. In the near-field case, the active RIS achieves $25.2858~\text{bps/Hz}$, while the passive RIS reaches only $13.7580~\text{bps/Hz}$. This gap becomes even more pronounced in the far-field scenario, where the active RIS attains $24.5325~\text{bps/Hz}$, nearly three times higher than the passive RIS’s $8.6790~\text{bps/Hz}$, under identical pilot configurations. These results clearly indicate that the active RIS not only enables more accurate channel estimation but also substantially enhances the achievable data rate of the overall communication system.
A key observation is, when the user is located in the far field, both approximation models provide accurate results. However, in the near field, the far-field approximation becomes less effective. The performance gap in the active RIS case remains minimal— a difference of $1.7392~\text{bps/Hz}$—whereas for the passive RIS case, it widens significantly to $5.2515~\text{bps/Hz}$, reaffirming the superiority of the active RIS architecture.
@@
For any fixed number of pilots, the achievable rate of the active RIS is consistently and markedly higher than that of its passive counterpart. This highlights the pilot efficiency of the active RIS: it requires considerably fewer training resources to achieve near-optimal performance, thereby reserving more channel uses for data transmission and improving overall spectral efficiency.
These findings validate the efficacy of employing active RIS for channel estimation. Importantly, this performance gain is realized under a fair total power budget, with the active RIS system consuming less transmit power at the base station. These results demonstrate that allocating a fraction of the system’s power budget to the RIS for signal amplification is far more beneficial than concentrating all power at the transmitter in a passive setup. By effectively overcoming the double-path-loss effect, the active RIS emerges not merely as a viable alternative but as a superior paradigm for achieving high-reliability, high-efficiency wireless communication.
@@ VI Conclusion
Giving every RIS element its own amplifier fundamentally changes the trade-offs relative to passive designs. By converting multiplicative path loss into net link gain, active RIS can increase capacity and efficiency in 6G networks. The future of the wireless network depends on practical scenarios of hardware implementation, such as the nonlinearity of amplifiers, tight power budgets, and calibration drift. Some semi-RIS architectures, where some elements act as active and others as passive, are also worth exploring by using sparse sensing with adaptive gain control to scale wideband estimation. For the future, the most important thing should be building hybrid approaches that focus not only on the bending of model-based reasoning but also on data-driven learning approaches.  It will then be easily adaptable to the user’s mobility and in cases of channel’s blockage while remaining within the limits of thermal and power constraints. If the designs of the codebooks, algorithms, and hardware are integrated, they will deliver many meaningful results in terms of both spectral and energy efficiency for the upcoming rise of the active RIS in the next generation of wireless networks.

### v2

@@ IV Adaptive Beam and Amplification Factor
3:for $\ell=2$ to $L$ do
4:  Measurement: $\mathbf{y}_{\ell}\leftarrow\sqrt{P_{p}}\,\mathbf{B}_{\ell}\mathbf{D}_{h}\mathbf{g}+\mathbf{v}_{\ell}+\mathbf{w}_{\ell}$
5:  where $\mathbf{v}_{\ell}(i)=\mathbf{B}_{\ell}(i,:)\mathbf{D}_{h}\mathbf{v}_{i}$
6:  Parametric MLE:
7:  $\hat{\bm{\psi}}_{\ell}=\arg\max_{\bm{\psi}\in\Psi}\frac{\big|\mathbf{y}_{\ell}^{H}\mathbf{F}^{-1}\mathbf{B}_{\ell}\mathbf{D}_{h}\bm{a}(\bm{\psi})\big|^{2}}{\bm{a}(\bm{\psi})^{H}\mathbf{D}_{h}^{H}\mathbf{B}_{\ell}^{H}\mathbf{F}^{-1}\mathbf{B}_{\ell}\mathbf{D}_{h}\bm{a}(\bm{\psi})}$
8:  Closed-form updates: compute $\hat{\beta}_{\ell}$, $\hat{\omega}_{\ell}$ and set $\hat{\mathbf{g}}_{\ell}$
9:  Active amplitude shaping:
10:  for $k=1$ to NN do
11:   $p_{k}^{(\mathrm{est})}\leftarrow C\,\alpha_{k}/(\beta_{k}+\gamma_{k})$
12:  end for
13:  if $\ell<L$ then
14:   Phase steering:
15:   $)\displaystyle\bar{\bm{\theta}}_{\ell}^{\rm opt}\leftarrow\exp\!\big(-j\,\angle(\mathbf{D}_{h}\hat{\mathbf{g}}_{\ell}$$\big)$,
16:   $\bm{\Phi}_{\ell}^{\star}\leftarrow\mathbf{p}^{(\mathrm{est})}\odot\bar{\bm{\theta}}_{\ell}^{\rm opt}$.
17:   Pilot selection: $\bm{\theta}_{\ell+1}\leftarrow\arg\max_{\bm{\theta}\in\bm{\Theta}}\big|(\bm{\Phi}_{\ell}^{\star})^{H}\bm{\theta}\big|$
18:   Append & update:
19:   $\bm{\Phi}_{\rm new}\leftarrow\mathbf{p}^{(\mathrm{est})}\odot\bm{\theta}_{\ell+1}$,
20:   $\mathbf{B}_{\ell+1}\leftarrow\begin{bmatrix}\mathbf{B}_{\ell}\\ \bm{\Phi}_{\rm new}^{\top}\end{bmatrix}$,
21:   $\bm{\Theta}\leftarrow\bm{\Theta}\setminus\{\bm{\theta}_{\ell+1}\}$.
22:   Transmit pilot $\ell{+}1$; acquire $\mathbf{y}_{\ell+1}$
23:  end if
24:end for
25:Output: $\hat{\mathbf{g}}\leftarrow\hat{\mathbf{g}}_{L}$
Figure 1: NMSE versus the pilot length when the user is at a random location in the near-field of the active RIS.
Figure 2: The average Achievable Rate versus the pilot length when the user is at a random location in the near-field of the active RIS.
Figure 3: NMSE versus the pilot length when the user is at a random location in the far-field of the active RIS.
Figure 4: The average Achievable Rate versus the pilot length when the user is at a random location in the far-field of the active RIS.
Figure 5: Achievable Rate versus the percentage of total power given to active RIS for three different total powers.
@@ V Simulation Results Analysis
We evaluate the proposed estimator via Monte Carlo simulations at a carrier of $28$ GHz with $1$ MHz bandwidth. The system has a single-antenna BS & UE assisted by an active RIS (ARIS): a $32\times 32$ UPA with half-wavelength spacing ($\Delta_{H}=\Delta_{V}=\lambda/2$) placed $15$ m from the BS. User angles follow $\phi,\varphi\sim\mathcal{U}[-\pi/3,\pi/3]$. Near vs. far field is decided by the Fraunhofer distance $d_{\mathrm{f}}=2D^{2}/\lambda$, where $D$ is the largest aperture dimension. For the near field, $r\sim\mathcal{U}[d_{\mathrm{B}},\,d_{\mathrm{f}}/10]$ with $d_{\mathrm{B}}=2D$ the Björnson distance. For the far field $r\sim\mathcal{U}[d_{\mathrm{f}},\,5d_{\mathrm{f}}]$. Pilot SNR is taken as $10$ dB higher than the data SNR.
We benchmark against a conventional passive RIS under a fixed total power budget of $200$ mW. For the passive RIS, the entire budget is allocated for the BS. For the active RIS, $25\%$ of the budget powers the RIS internal amplifiers, and the BS radiates the remaining $75\%$, ensuring a fair comparison. Results are averaged over $1000$ independent channel & noise realizations.
With the adaptive pilot strategy and wide beams initialization technique for the Active RIS, under both near and far field regimes, the curve converges in around 7- 8 pilots.
Figs. 4 & 4 plot capacity bounds and pilot-aided achievable rates. For any pilot count, the active RIS yields higher rates, and its operational rate closely approaches its capacity bound, whereas the passive RIS remains well below. In the near field, the active RIS achieves $25.29$ bps/Hz vs. $14.31$ bps/Hz for the passive RIS. In the far field, the multiplicative fading effect can be seen much more clearly, where the active RIS attains nearly sixfold $21.01$ bps/Hz in comparison to $3.87$ bps/Hz for the passive RIS, under identical pilot configurations.
@@
In the far field, 4 , the active RIS attains a final NMSE several dB below the minimum NMSE of the passive RIS, under both near and far-field approximation models. In the near field, 4 , the two architectures exhibit almost equivalent behavior under both models.
When the user is in the far field, both approximation models are nearly accurate. However, in the near field, the far-field approximation degrades. Due to not having any information regarding the distance, the far-field approximator can not reach its capacity.
@@
For any fixed number of pilots, the active RIS consistently and markedly outperforms the passive RIS, requiring fewer training resources to achieve near-optimal performance, thereby reserving more channel uses for data and improving spectral efficiency.
Figure 5 shows the Capacity versus the power-splitting factor $\alpha=P_{\mathrm{RIS}}/P_{\mathrm{tot}}$ for $P_{\mathrm{tot}}\in\{200,500,1000\}$ mW. Increasing $\alpha$ initially improves the rate due to stronger RIS amplification, but the gain saturates, and drops as $\alpha\!\to\!1$ since the user transmit power $(1-\alpha)P_{\mathrm{tot}}$ vanishes, yielding an optimal split at intermediate-to-high $\alpha$.
Active RIS achieves better channel estimation and higher rates under the same total power budget, even with lower BS transmit power, since allocating power to RIS amplification mitigates double-path loss, making active RIS preferable to passive designs for reliable, high-efficiency links.
@@ VI Conclusion
Equipping each RIS element with an amplifier fundamentally shifts the design trade-offs versus passive RIS: by turning multiplicative path loss into net link gain, active RIS can boost capacity and energy efficiency in 6G. Practical deployment, however, must confront hardware realities: amplifier nonlinearity, tight power budgets, and calibration drift. Semi-active designs that mix active and passive elements are also promising, e.g., sparse sensing with adaptive gain for scalable wideband estimation. Looking ahead, hybrid approaches that fuse model-based reasoning with data-driven learning can adapt to mobility and blockage while respecting thermal and power limits. Integrating codebooks, algorithms, and hardware co-design is likely to deliver meaningful spectral and energy efficiency gains as ARIS matures for next-generation networks.